A coil of the turn area S= 200cm2 is mounted in a uniform magnetic field B = 8.10-4 T. There is a current of 20 A in the coil, which has 25 turns. When the plane of the coil makes an angle of 600 with direction of the field, what is the torque tending to rotate the coil?


1
Expert's answer
2021-04-26T18:55:19-0400

To be given in question

Coil area(A) =200"cm^2"

Coil magnetic field(B)="8\times10^{-4} T"

Current(i)=20Amp

Turns(n)=25

"\theta=60°"

To be asked in question

"\tau =?"

We know that

"M=niA\rightarrow(1)"

"\tau=M\times B"

"\tau=MBsin\theta\rightarrow(2)"

eqution (1)and(2) we can written as

"\tau=niABsin\theta\rightarrow(3)"

eqution (3)put values


"\tau=25\times20\times200\times10^{-4}\times8\times10^{-4}\times sin60°" "\tau=4\sqrt{3}\times10^{-3}Nmeter"

"\tau=6.9282\times10^{-3} Nmeter"